# A General Bayesian Framework for Informative Input Design in System Identification

### v1

@@ Abstract
We tackle the problem of informative input design for system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data. We propose a methodology that is compatible with any system and parametric family of models. Our approach only requires input-output data from the system and first-order information from the model with respect to the parameters. Our algorithm consists of two modules. First, we formulate the problem of system identification from a Bayesian perspective and propose an approximate iterative method to optimize the model’s parameters. Based on this Bayesian formulation, we are able to define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input. Our method outperforms model-free baselines with various linear and nonlinear dynamics.
@@ I Introduction
Controlling uncertainty is a central problem for various engineering fields. It abstracts to the selection of inputs to a system in order to reduce an uncertainty measure for a variable of interest. In robotics, the problem usually coincides with adaptive informative input design, where trajectories are planned to maximize the information gathered from the unknown environment [1]. Similarly, robots may choose actions to reduce the uncertainty in their state [2]. In systems engineering, we may want to choose inputs to a system to improve the estimate of system dynamics [3].
@@
In this work, we are interested in the problem of informative input design for the modeling of a system. Our formulation tackles the general problem of dataset collection for model-fitting purposes. Ott et al. have shown that selecting control inputs based on the current model can improve the system identification process [4]. Informative input design can reduce the data gathering requirements for the purpose of modeling the system, which can be costly in real-world applications. This can be particularly useful in high-dimensional problems, where finding informative directions to perturb the system can be challenging. Informative input design can also be regarded as the first step in a more complicated pipeline that involves the control of the underlying system, e.g., with model predictive control [5]. The problem of selecting input sequences that lead to highly informative measurements under constraints is still an open problem [4].
@@
There are four basic elements in the problem of informative input design. We need to select a parametric family of models for the underlying system. Our parameter estimate is improved once more data becomes available and we must choose the optimization approach to update the model parameters. The model family must be expressive enough to capture the true system, but also simple enough to allow the optimization algorithm to converge in a meaningful manner. Furthermore, we need an uncertainty measure to minimize over the next input. This allows us to select inputs that best inform our choice of model parameters. The last minimization problem must also be accompanied by an optimization approach.
@@
These specifications make designing methods for informative input design challenging and lead to algorithms that are usually limited to a specific set of parametric models. Our approach allows for the selection of any parametric family of models for the true system, although we do not focus on the suitability of the chosen model. It leverages first-order information for the model and successive convex problems based on the first-order Taylor expansion to update the model parameters. This simplifies the model update step discussed in the previous paragraph. Because the first-order Taylor expansion gives rise to a linear Gaussian model, we use an uncertainty measure that is accurate in the linear Gaussian case for the input design.
@@
The rest of the paper is organized as follows. In section II we discuss prior work. In section III, we formulate the problem of informative input design, while in section IV we describe our proposed approach. We test our algorithm in various scenarios in section V and the paper concludes with section VI.
@@ II Prior Work
The problem of informative input design has been well studied. Early research on the optimal selection of inputs dates back to the 1960s [6, 7] and highlights the importance of the problem for the identification of systems. Recent advances in computational capabilities and learning-based methods have once again brought attention to the problem [8, 9]. This section summarizes three categories of informative input design: model-free methods, methods for linear systems, and methods for nonlinear systems.
Informative inputs can be designed by ignoring the current estimate of the system dynamics. These methods are called model-free because they are not based on a model of the system. One such method is the orthogonal optimized multisine input design [10], which has found applications in the study of aircraft stability and control. This technique produces inputs that are mutually orthogonal in both the time and frequency domains, in an attempt to perturb the system in orthogonal directions. Similarly, pseudorandom binary sequences can be used to excite the system [11]. The idea is to leverage the sequence’s randomness in order to excite the system in interesting ways. Our method differs from model-free approaches because it uses the currently estimated model to inform the next input design.
Although they are simple, linear models are a powerful representation tool. They usually allow for closed-form analysis and are easily interpretable from a statistical standpoint. Informative input design using such models has a vast literature. A recent method that aims to estimate linear dynamics suggests selecting the next input by minimizing the trace of the covariance of the model parameters with linear dynamics constraints [4]. By linearizing the objective, the authors obtain a convex problem that is easily solvable for the next input. This method is compatible with the dynamic mode decomposition with control (DMDc) framework that produces reduced order models for the dynamics [12]. Our method can provide an approximate solution to any parametric model family and is exact in the case of linear models.
@@
In various applications, the dynamics contain modes that are not properly described by a linear model. In these cases, we must design inputs for nonlinear system identification. Early results focused on specific dynamics for which the input design problem can either be relaxed using linear matrix inequalities [13] or techniques from linear systems can be applied to design informative step inputs [14]. In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification. In this case, convex-based methods exist for informative input design [15]. In contrast with these methods, our proposed algorithm is applicable in the case of general nonlinear models.
@@
Goodwin presented an algorithm for informative input design in the general nonlinear case that is based on linearization of the parametric model [16]. Our work adds to this work in an important way: we present a method to update the estimate of the model parameters that respects the linearization and is based on convex optimization, which makes it computationally fast. A similar work was done by Babar and Baglietto [17], who further propose an extended Kalman filter for the parameter estimation. We improve upon their work by adding a trust-region constraint to the update of the model parameters after the linearization and by proposing a modified optimization objective for the selection of the next input. Our objective depends on a parameter that directly encourages the satisfaction of any constraints, i.e., dynamics constraints. In contrast with both works [16, 17], we present a fine-tuning method for the assumed noise covariances that can improve convergence.
@@ III Problem Formulation
where $x\in\mathbb{R}^{d_{x}},y\in\mathbb{R}^{d_{y}}$ and $g:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$. If $g$ represents dynamics, the input is usually the current state and control, which we abstract as $x$, while the next state is abstracted as $y$. We refer to $x$ as the input and $y$ as the output.
@@
We opt for a Bayesian treatment of the problem. We suppose that the inputs and outputs are exactly observed, because our main goal is efficient dataset acquisition. Our method easily extends to the case of noisy observation of the outputs.
@@ IV Proposed Method
We present our methodology for the design of an informative input in system identification. We assume that we have already observed input-output data points $(x_{i},y_{i})_{i=1}^{n}$ from the dynamics:
@@
We first outline our Bayesian approach for improving our current estimate of the model parameters. We then use this Bayesian formulation in order to define an optimization problem for the next input. Finally, we discuss some fine-tuning that can improve performance before bringing the concepts together in our proposed algorithm.
@@ IV-A Parameter Estimation
The collected data satisfy eq. 2. Because the current model $f(\cdot;\theta)$ might not match the true system $g(\cdot)$ exactly, it holds that
$y_{i}=f(x_{i};\theta)+\epsilon_{i},$ (3)
for some noise realization $\epsilon_{i}$. We choose the underlying noise model $\epsilon_{i}\sim\mathcal{N}(0,\Sigma_{i})$. Assuming conditionally independent measurements, the posterior for the hidden parameter is given by
@@
where we choose the prior $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$. The noise and prior are assumed Gaussian because they lead to a convex maximum likelihood problem [18].
The maximum a posteriori estimate for the parameter under these assumptions is given by
@@
$\qquad\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma_{i}^{-1}}^{2}.$
In general, this can be a very difficult problem to solve because it depends on the structure of the parametric family $f$ with respect to the parameter $\theta$. We opt for an approximate treatment of the optimization problem that lends itself to a convex solver.
Suppose that the current parameter estimate is $\hat{\theta}$. We do not allow our next estimate to differ much relative to the current one. This is justified because the next input-output pair should not change our estimate drastically. For model parameters close to $\hat{\theta}$, the parametric family can be accurately approximated using a first order Taylor series. We write
@@
Plugging (6) into (5) and imposing the distance constraint, we get that
$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta\lVert\hat{\theta}\rVert}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}$ (8)
$+\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x;\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma_{i}^{-1}}^{2}.$
This is a convex problem with respect to $\theta$ and can easily be handled by a computational framework for convex optimization [19]. The constraint guarantees that we search for the next model parameter in the region where the current Taylor series is accurate. It corresponds to a trust region.
@@ IV-B Informative Input Design
$\Sigma_{\mathrm{posterior}}^{-1}=\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+}).$ (9)
It has been shown that such an expression can be accurate in the nonlinear case, although it is exact only in the linear regime [20].
@@
$\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+})\right),$ (10)
@@
$\operatorname*{arg\;max}_{x_{n+1}}$ $\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+})\right)$ (11)
@@
Equation 9 defines the posterior covariance for the parameter $\theta$. At the next iteration , we can replace the constraint of eq. 8 with
$\lVert\theta-\hat{\theta}\rVert^{2}_{\Sigma_{\mathrm{posterior}}^{-1}}\leq\chi_{d,p}^{2},$
where $\chi_{d,p}^{2}$ is the value of the chi-squared distribution of $d$ degrees of freedom at the $p$-percentile. This would allow $\theta$ to vary within the contours that correspond up to the $p$-th percentile of the posterior Gaussian distribution. Choosing the adequate constraint in eq. 8 involves balancing this statistical interpretation with the accuracy of the Taylor series.
IV-C Updating the Covariances $\Sigma_{i}$
@@
Using these samples, we can compute a sample covariance for the overall error in our model
$\hat{\Sigma}_{+}=w\Sigma_{\mathrm{meas,prior}}+\dfrac{1-w}{n}\sum_{t=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top},\ w\in\left[0,1\right],$ (13)
i.e., a convex combination of the empirical covariance and a prior covariance. This can be utilized in later iterations.

### v2

# An Iterative Bayesian Approach for System Identification based on Linear Gaussian Models

@@ Abstract
We tackle the problem of system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data. We propose a flexible and computationally tractable methodology that is compatible with any system and parametric family of models. Our approach only requires input-output data from the system and first-order information from the model with respect to the parameters. Our algorithm consists of two modules. First, we formulate the problem of system identification from a Bayesian perspective and use a linear Gaussian model approximation to iteratively optimize the model’s parameters. In each iteration, we propose to use the input-output data to tune the covariance of the linear Gaussian model. This statistically calibrates the approach. Secondly, we define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input. We test our method with linear and nonlinear dynamics.
@@ I Introduction
In this work, we are interested in the problem of fitting a model to a system and choosing informative inputs to improve the model. Our formulation tackles the general problem of model-fitting and dataset collection, which is known as system identification. Obtaining a good model of the system is a key requirement in various engineering tasks. System identification is usually the first step in a larger pipeline that involves the control of the underlying system, e.g., with model predictive control [1]. Fitting model parameters to data can be challenging when the model is a complicated function of the parameters. In addition, Ott et al. have shown that selecting control inputs based on the current model can improve the system identification process [2]. Informative input design can reduce the data gathering requirements for the purpose of modeling the system, which can be costly in real-world applications. This can be particularly useful in high-dimensional problems, where finding informative directions to perturb the system can be challenging.
@@
There are four basic elements in the problem of system identification. We need to select a parametric family of models for the underlying system. The family must be expressive enough to capture the true system, but also simple enough to optimize over. Our parameter estimate is improved once more data becomes available and we must also choose the optimization approach that updates the model parameters. A vast literature exists on methods for estimating the model parameters [3]. Furthermore, we need an uncertainty measure to minimize over the next input. This allows us to select inputs that best inform our choice of model parameters. Various objectives based on the Fisher information matrix can be minimized over the input [4]. This optimization problem must also be accompanied by an optimization approach.
The specifications discussed above make designing methods for system identification challenging and lead to algorithms that are usually limited to a specific set of parametric models. Our work deals with the following question:
  • Can we design a method for system identification that is computationally tractable, compatible with any parametric model family and able to suggest whether a parametric model family is suitable for a given system?
@@
Our approach leverages first-order information to construct a linear Gaussian model for the output given the input, as a function of the model parameters. Based on the first-order Taylor expansion, we solve successive convex problems, corresponding to maximum likelihood estimation (MLE), to update the model parameters. This simplifies the model update step discussed in the previous paragraph, which becomes computationally tractable. We also propose a procedure to tune the covariance associated with the linear Gaussian model. This allows us to statistically calibrate our model and judge the suitability of the parametric model family. Finally, using the linear Gaussian model from the first-order Taylor expansion, we define an uncertainty measure that is accurate in the linear Gaussian case for the input design.
@@
The rest of the paper is organized as follows. In section II we discuss prior work. In section III, we formulate the problem of system identification, while in section IV we describe our proposed approach. We test our algorithm in various scenarios in section V and the paper concludes with section VI.
@@ II Prior Work
The problem of system identification has been well studied. Early research on the optimal selection of inputs dates back to the 1960s [5, 6] and highlights the importance of the problem for the identification of systems. Recent advances in computational capabilities and learning-based methods have once again brought attention to the problem [7, 8]. This section summarizes the prior work on parameter estimation for system identification and informative input design.
Estimating the model parameters using input-output data is usually formulated as a regression problem [9]. A regression problem can be equivalently viewed as MLE under an appropriate distribution. For example, least squares regression corresponds to MLE with a Gaussian distribution. When estimating the model parameters, it is common to assume the true system belongs to the parametric model family [3, 4, 10]. Then, under regularity conditions on the distribution of the input or the structure of the system, the parameter estimate is concentrated around the true system parameter [3]. Our work deals with the problem of system identification when the true system does not necessarily belong to the selected parametric model family. In this case, it is critical to obtain a good fit but also a sense of the suitability of the selected model family.
Under the assumptions mentioned in the previous paragraph, we usually obtain an unbiased estimate for the model parameters. This implies that we can design informative inputs by controlling the covariance of the estimator. Assuming an efficient estimator, the covariance converges to the inverse Fisher information matrix [4]. Therefore, a scalar measure of the Fisher information matrix is usually optimized over the inputs. Different scalar functions of the Fisher information matrix are used. For example, a D-optimal input maximizes the determinant of the Fisher information matrix [11]. This corresponds to minimizing the uncertainty volume of the parameter estimate. Our approach optimizes any desired scalar function of an approximate Fisher information matrix to determine the next input. Usually, the informative input design is done before the parameter estimation, but active techniques also exist [12]. Our approach is a simple and generally applicable active system identification approach.
@@
In various applications, the system contains modes that are not properly described by a linear model. In these cases, we must deal with nonlinear models and nonlinear system identification. Early results focused on specific system dynamics for which the input design problem can either be relaxed using linear matrix inequalities [13] or techniques from linear systems can be applied to design informative step inputs [14]. In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification [4]. There is consensus among researchers that nonlinear system identification is much more difficult than linear system identification [15]. A simple justification can be made by examining the difficulty of solving a nonlinear regression problem. The problem of informative input design for nonlinear dynamical systems remains a large unexplored area of research [4]. Our approach deals with the identification of nonlinear dynamics.
@@
We examine methods for nonlinear system identification that are relevant to our approach. Goodwin presented an algorithm for informative input design in the general nonlinear case that is based on linearization of the parametric model [16]. Our work adds to this work in an important way: our method updates the estimate of the model parameters in a way that respects the linearization and is based on convex optimization, which makes it computationally fast. A similar work was done by Babar and Baglietto [17], who propose an extended Kalman filter for the joint state and parameter estimation of the dynamical system. Model predictive control techniques are used to select the next input. We improve upon their work by adding a trust-region constraint to the update of the model parameters after the linearization and by proposing a modified optimization objective for the selection of the next input that does not rely on sampling. In contrast with both works [16, 17], we tune the noise covariance at every iteration. This statistically calibrates our approach and provides information on the suitability of the selected model family for the system in question. We show how the approach can fail without tuning the covariance.
@@ III Problem Formulation
where $x\in\mathbb{R}^{d_{x}},y\in\mathbb{R}^{d_{y}}$ and $g:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$. If $g$ represents dynamics, the input is usually the current state and control, which we abstract as $x$, while the next state is abstracted as $y$. We refer to $x$ as the input and $y$ as the output. We treat system $g$ as a black-box. We can only supply an input $x$ and provide an output $y$.
@@ IV Proposed Method
We present our methodology for system identification. We opt for a Bayesian treatment of the problem. We suppose that the inputs and outputs are exactly observed. Our method easily extends to the case of noisy observation of the outputs. We assume that we have already observed input-output data points $\{(x_{i},y_{i})\}_{i=1}^{n}$ from the system:
@@
We first outline our Bayesian approach for improving the current estimate of the model parameters. We then use this Bayesian formulation in order to define an optimization problem for the next input. The components presented so far can be found in the literature [16], with the exception of the trust region in the estimate improvement step and the ability to choose the noise probability density function. Finally, we discuss the fine-tuning, which is critical in the convergence of the method and constitutes our major contribution.
@@ IV-A Parameter Estimation
The collected data satisfy eq. 2. Because the current model $f(\cdot;\hat{\theta})$ might not match the true system $g(\cdot)$ exactly, it holds that
$y_{i}=f(x_{i};\hat{\theta})+\epsilon_{i},$ (3)
for some noise realization $\epsilon_{i}$. We choose the underlying noise model as $\epsilon_{i}\sim\mathcal{N}(0,\Sigma)$ for some covariance $\Sigma$. Tuning this covariance is critical in the performance of the algorithm. Assuming conditionally independent measurements, the posterior for the hidden parameter is given by
@@
where we choose the prior $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$.
The maximum a posteriori (MAP) estimate for the parameter under these assumptions is given by
@@
$\qquad\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma^{-1}}^{2}.$
In general, this can be a very difficult problem to solve because it depends on the structure of the parametric family $f$ with respect to the parameter $\theta$. We opt for an approximate treatment of the optimization problem that lends itself to a convex solver [18].
Suppose that the current parameter estimate is $\hat{\theta}$. For model parameters close to $\hat{\theta}$, the parametric family can be accurately approximated using a first order Taylor series. We write
@@
Plugging (6) into (5) and imposing a trust region constraint, we get that
$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}$ (8)
$+\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x_{i};\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma^{-1}}^{2}.$
This is a convex problem with respect to $\theta$ and can easily be handled by a computational framework for convex optimization [19]. The constraint $\lVert\theta-\hat{\theta}\rVert\leq\delta$ guarantees that we search for the next model parameter in the region where the Taylor series for $f(\cdot;\theta)$ is sufficiently close to $f(\cdot;\theta)$. It also does not allow our next estimate to differ much relative to the current one. This is justified because the next input-output pair should not change our estimate drastically. Placing too much trust on any one datapoint can lead to excessively large steps or premature convergence [20].
We can replace Gaussian densities with any other log concave density and still obtain a convex problem. The used density for $\epsilon_{i}$ can be changed in each estimation step, based on the realized errors $\{y_{i}-f(x_{i};\hat{\theta})\}_{i=1}^{n}$. This allows us to select among a family of parameter estimates. More details on this will be given in section IV-C.
@@ IV-B Informative Input Design
$\Sigma_{\mathrm{posterior}}^{-1}=\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+}).$ (9)
It has been shown that such an expression for the covariance can be accurate in the nonlinear case, although it is exact only in the linear regime [21]. Eq. (9) is an approximate Fisher information matrix, where a prior is added and the expectation over the data is not used. We note that it only depends on the inputs $x_{1},\dots,x_{n+1}$ and not on the corresponding outputs.
@@
$\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+})\right),$ (10)
@@
$\operatorname*{arg\;max}_{x_{n+1}}$ $\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+})\right)$ (11)
@@
IV-C Finetuning
We present the finetuning approach for the covariance $\Sigma$ that is critical for the convergence of our algorithm. We also propose a way to tune $\delta$.
@@ IV-C Finetuning
Using these samples, we can compute a covariance for the errors caused by the mismatch between $g(\cdot)$ and $f(\cdot;\hat{\theta}_{+})$:
$\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top}.$ (13)
However, in problem (8) we consider the distribution of the errors
$y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\theta$ (14)
that consist of two terms: the error from the mismatch between $f(\cdot;\theta)$ and $g(\cdot)$ and the error because of the linearization, since
$y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\hat{\theta}_{+}=$ (15)
$\hat{\epsilon}_{i}+\underbrace{f(x_{i};\hat{\theta}_{+})-\left(b(x_{i};\hat{\theta})+C(x_{i};\hat{\theta})\hat{\theta}_{+}\right)}_{\bar{\epsilon}_{i}}.$
This implies that the covariance $\Sigma$ should be set using the sample covariance of $\{\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\}_{i=1}^{n}$:
$\Sigma\leftarrow\dfrac{1}{n}\sum_{i=1}^{n}\left(\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\right)\left(\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\right)^{\top},$ (16)
where $\bar{\epsilon}_{i}$ is the error between the value from the model and the value from the Taylor expansion for the $i$-th output.
For setting the trust region $\delta$, we can use a simple heuristic. If the average linearization error size $n^{-1}\sum_{i=1}^{n}\lVert\bar{\epsilon}_{i}\rVert$ becomes comparable to the average absolute model error size $n^{-1}\sum_{i=1}^{n}\lVert\hat{\epsilon}_{i}\rVert$, we reduce $\delta$. This procedure guarantees that we keep any error due to linearization small compared to the errors because of the model mismatch. This allows for confident steps in the parameter update.